# Learning Plasma Dynamics and Robust Rampdown Trajectories with Predict-First Experiments at TCV

### v1

@@ Abstract
The rampdown in tokamak operations is a difficult to simulate phase during which the plasma is often pushed towards multiple instability limits. To address this challenge, and reduce the risk of disrupting operations, we leverage recent advances in Scientific Machine Learning (SciML) to develop a neural state-space model (NSSM) that predicts plasma dynamics during Tokamak à Configuration Variable (TCV) rampdowns. By integrating simple physics structure and data-driven models, the NSSM efficiently learns plasma dynamics during the rampdown from a modest dataset of 311 pulses with only five pulses in the reactor relevant high performance regime. The NSSM is parallelized across uncertainties, and reinforcement learning (RL) is applied to design trajectories that avoid multiple instability limits with high probability. Experiments at TCV ramping down high performance plasmas show statistically significant improvements in current and energy at plasma termination, with improvements in speed through continuous re-training. A predict-first experiment, increasing plasma current by 20% from baseline, demonstrates the NSSM’s ability to make small extrapolations with sufficient accuracy to design trajectories that successfully terminate the pulse. The developed approach paves the way for designing tokamak controls with robustness to considerable uncertainty, and demonstrates the relevance of the SciML approach to learning plasma dynamics for rapidly developing robust trajectories and controls during the incremental campaigns of upcoming burning plasma tokamaks.
1 Introduction
Upcoming burning plasma tokamaks, such as SPARC [2] and ITER [3], require reliable plasma control to avoid operational delays and machine damage due to plasma disruptions, a challenge that will only increase for tokamak pilot plants [4] like ARC [5] and DEMO [6]. Given this risk becomes intolerable at high plasma current, $I_{p}$, and stored energy, $W_{tot}$, a key mitigation strategy is to de-energize the plasma by performing a rampdown of the plasma current, but doing so typically pushes the plasma closer to multiple instability boundaries [7, 8, 9]. Figure 1 depicts the phases of a tokamak pulse, beginning with rampup of the plasma current to the steady-state flattop phase, and ending with a rampdown. Notably, Figure 1 also shows an example of a quantity correlated with plasma instability growing during the rampdown phase, a challenge which motivates the algorithmic design of safe rampdown trajectories. This challenge is especially acute in reactor relevant high performance plasmas, which operate close to instability boundaries to achieve the high normalized plasma density, typically quantified by the Greenwald fraction $f_{GW}$ [10], and normalized plasma pressure, $\beta_{N}$ [11], necessary for economical energy production. The importance of designing robust rampdowns for reactor relevant fusion plasmas is highlighted by the recent record-breaking high performance campaign at the Joint European Torus (JET), for which most disruptions occurred during the termination phase [12]. For the baseline scenario, a $\approx 15\%$ increase of the plasma current, from 3MA to 3.5MA, increased the disruptivity considerably from $\approx 20\%$ to $\approx 50\%$ [12]. This challenge motivates the development of tools that can rapidly adapt rampdown trajectories to manage disruptivity as fusion performance is increased.
@@ 1 Introduction
The RL environment is massively parallelized to design trajectories with robustness to uncertainties, including the initial conditions of the plasma and its time varying dynamics. In contrast with prior works on applying RL approaches to fusion, we leverage its capabilities for offline design of robust trajectories, which is more readily applicable to the safety-critical settings of upcoming tokamaks than RL for real-time control, as was done in prior works [29, 30]. After a small number of initial trial shots, the plasma reliably terminated at low plasma current and stored energy for five consecutive high performance shots, with statistically significant improvements relative to baseline, although we encourage caution in interpreting the statistics of the result due to the small sample size. As a test of the viability of this approach for performing small extrapolations in an incremental high performance campaign, which upcoming tokamaks like SPARC and ITER will undergo, we design a rampdown trajectory and perform a predict-first experiment by increasing the plasma current by 20%, from 140kA to 170kA, for a high $\beta_{N}$ plasma near the density limit, a scenario for which zero shots of rampdown data exists for TCV. In this extrapolation test, we a priori predict the dynamics of key quantities to within sufficient accuracy to successfully terminate the plasma on both attempts.
@@
Figure 1: Data from an example TCV pulse (#73919), showing the rampup, flattop, and rampdown phases, which are defined by the plasma current $I_{p}$. The bottom subplot shows how the rampdown pushes the plasma closer to an instability limit, in this case the Greenwald density limit, defined by the Greenwald fraction $f_{GW}=1$. Note this limit is approximate due to the, at present, incomplete physics understanding of the density limit [31, 32]. Also note that the flattop phase is abbreviated here to more clearly highlight the rampup and rampdown.
2 Results
2.1 Experiment Overview
@@
The reported experiment was conducted as a part of the 2024 TCV integrated control, high performance experimental campaign. Flattop plasmas operated at a high performance of $\beta_{N}>2.0$ and near the density limit with a highly elongated diverted geometry with $\kappa\approx 1.6$ and $q_{95}\approx 4$. Initial shots in the experiment operated at $I_{p}=140$ kA, henceforth known as the baseline high performance (HP) scenario, with a final extrapolation test at $I_{p}=170$ kA. Successful rampdowns from these scenarios require careful management of multiple plasma instability limits that can be exacerbated by details of the plasma trajectories. To name a few considerations, an overly fast reduction in plasma current increases the Greenwald fraction, prompting a density limit event, managing the relatively high elongation diverted geometry introduces risks of a vertical displacement event (VDE), and fast reductions in elongation and minor radius can decrease the safety factor $q_{95}$, exciting magnetohydrodynamic (MHD) instabilities. Managing the density limit in this scenario is a particular challenge for fast terminations, as the relatively long particle confinement time-scale is a major constraint on the speed of the rampdown.
@@
To address the problem, a NSSM dynamics model was trained on a modest dataset of past rampdowns, which contains only 5 incomplete rampdowns in the relevant high performance parameter space, as shown in Figure 2A-B. This NSSM is then used in a reinforcement learning (RL) environment to optimize a reward function, designed to minimize time to a goal plasma current of $40$kA and stored energy of $0.5$kJ without disrupting, as shown in Figure 2C. The action space was chosen to be plasma current, $I_{p}$, shaping parameters $\kappa$ and $a_{minor}$ and neutral beam injection (NBI) power $P_{NBI}$. User-specified constraints were set on the Greenwald fraction $f_{GW}$, safety factor $q_{95}$, vertical instability growth rate $\gamma_{vgr}$, and poloidal beta $\beta_{p}$. The optimized action trajectories were then manually programmed into the TCV plasma control system (PCS). The details of the reward function, chosen limits, and PCS programming process are further discussed in the Methods.
@@
Every shot involved in the experiment is shown in Figure 2D, showing improvements on the normalized $I_{p}$ and $W_{tot}$ at time of plasma disruption over the course of the experimental runs. The un-optimized baseline rampdown trajectory for this scenario disrupted at relatively high current and stored energy in #81101 and #81102 at $I_{p}\approx 80$kA and $W_{tot}\approx 4$kJ. The experiment proceeded iteratively, with re-training of the NSSM on new data and trajectories done after shots #81635, #81745, #81751, and #81830. A preliminary optimized trajectory was deployed in TCV #81635, which reached the goal $I_{p}$ and $W_{tot}$ before disrupting, but post shot analysis showed poor radial control and tracking of the target shape, which was determined to be due to a legacy software issue detailed in Figure 14 in the Supplementary Information. Shots #81741 and #81745 were spent resolving this issue, with it properly resolved in #81751, as shown in Figure 14. #81751 still disrupted due to a VDE, which was found to be due to a large sensitivity of $\gamma_{vgr}$ to small control errors in the inner gap. After #81751, an uncertainty distribution on gap errors was added to the RL training environment to gain robustness to this uncertainty, with subsequent shots experiencing similar control errors but without similar increases in $\gamma_{vgr}$, demonstrating the importance of designing trajectories with robustness to real-world uncertainties, as further discussed in Section 2.3.
@@
Statistical significance of control results in fusion is typically difficult to establish due to the scarcity of experimental time and relevant data-points. This experiment also faces this, given the rampdown experiment involved only nine shots, two of which were dedicated to debugging a legacy software issue, with only five rampdowns in the database near the relevant high performance regime with $\beta_{N}>2$. We use these five shots as our control set and define two test sets: one with the debugging shots and one without. As shown in Figure 3, the Mann-Whitney U test [33] shows a statistically significant improvement in $W_{tot}$ , ($p<0.05$), at time of plasma termination of the experimental rampdowns for both definitions of the test set. Improvements in $I_{p}$ are only statistically significant when we do not include the debugging shots. While the results of this statistical test are encouraging, we urge caution in its interpretation given the small sample sizes involved, and the fact that tokamaks are highly drifting distributions in practice, with uncontrolled variables such as wall conditioning making a meaningful impact on experimental outcomes.
Figure 2: A) The training data distribution, with a modest dataset size at low performance and very few shots in the relevant high performance regime. The target scenarios for this work at 140kA and 170kA high normalized performance are shown. B) Depiction of the dynamics model training method, which involves comparing results from forward simulation of a NSSM against experimental data to compute the gradient of loss with respect to model parameters. C) Depiction of the trajectory optimization process. In addition to the trained dynamics model, the reinforcement learning (RL) training environment is defined by a reward function specifying the desired goal and a set of random variables that training environments are parallelized against, to find a trajectory that has robustness to uncertainties and off-normal events. D) The plasma current $I_{p}$ and stored energy $W_{tot}$ at plasma termination, relative to their values at the beginning of the termination phase for every shot during the experiment that reached the termination phase. E) A comparison of the current and stored energy trajectories for a baseline shot and an optimized shot, showing a faster, and non-disruptive, decrease in the plasma current and stored energy relative to the baseline.
$I_{p}$ $W_{tot}$
Test Only 0.024 0.0025
Test + Debug 0.12 0.0020
Figure 3: (Top) histograms of $I_{p}$ and $W_{tot}$ at time of plasma termination. (Bottom) p-values from the Mann-Whitney U test comparing performance of experimental shots relative to the control set of all shots in the database with $\beta_{N}$ > 1.5.
2.2 Medium-Scale Validation of NSSM Predictions
@@
The model underwent two training phases: an initial training phase on a wider dataset with 311 shots in the training dataset and 131 shots in the validation dataset. To improve the predictive power of the model for the relevant scenario, we fine-tune just the confinement scaling of the model by training only on 44 shots with $I_{p}\leq 200$kA, with all other model weights frozen. Due to the relatively small size of the fine-tuning dataset, we did not separate out a validation or testing dataset for this fine-tuning phase. As shown in Figure 4, the trained model is able to predict the time-dependent dynamics of key 0D kinetic and disruptive quantities to within tens of percent for full rampdowns in the validation dataset, even in the 95th percentile of error. The percent errors for $\gamma_{vgr}$ can be relatively large, but, as shown in Figure 15 this is largely attributable to the small value of $\gamma_{vgr}$ of limited plasmas as the absolute error is relatively low.
@@
The NSSM was initially developed with a neural network predictor for the kinetic profiles on the full $\rho$ grid, and initial training runs found that the profile predictor can accurately predict kinetic profiles given the set of 0D scalars listed in Table 2. Figure 5 provides an example comparison of predictions of the $T_{e}$ and $n_{e}$ profiles against Thomson measurements for a full shot in the validation dataset, showing accurate prediction across all phases of the shot. This result corroborates previous findings at NSTX-U that neural networks can accurately predict kinetic profiles given a set of similar 0D scalars [34]. Given this result suggests most of the relevant profile information is implicitly captured by 0D scalars, the profile predictor was disabled prior to running experiments to accelerate training, hence reported predictions of kinetic profiles are not predict-first. A noteworthy feature of this profile predictor that should be explored in future work is its ability to function as a Thomson up-sampler, as the input variables are all sampled at a higher time resolution, 1ms, than the TCV Thomson Scattering system, which takes measurements every 17ms.

### v2

@@ Abstract
The rampdown phase of a tokamak pulse is difficult to simulate and often exacerbates multiple plasma instabilities. To reduce the risk of disrupting operations, we leverage advances in Scientific Machine Learning (SciML) to combine physics with data-driven models, developing a neural state-space model (NSSM) that predicts plasma dynamics during Tokamak à Configuration Variable (TCV) rampdowns. The NSSM efficiently learns dynamics from a modest dataset of 311 pulses with only five pulses in a reactor-relevant high-performance regime. The NSSM is parallelized across uncertainties, and reinforcement learning (RL) is applied to design trajectories that avoid instability limits. High-performance experiments at TCV show statistically significant improvements in relevant metrics. A predict-first experiment, increasing plasma current by 20% from baseline, demonstrates the NSSM’s ability to make small extrapolations. The developed approach paves the way for designing tokamak controls with robustness to considerable uncertainty and demonstrates the relevance of SciML for fusion experiments.
Introduction
[…]
@@ Introduction
The RL environment is massively parallelized to design trajectories with robustness to uncertainties, including the initial conditions of the plasma and its time varying dynamics. We leverage the capabilities of RL for offline design of robust trajectories, which is more readily applicable to the safety-critical settings of upcoming tokamaks than RL for real-time control [29, 30, 31, 32]. A similar approach has previously been demonstrated at KSTAR for designing feed-forward trajectories that reach target states [33, 34]. After a small number of initial trial shots, the plasma reliably terminated at low plasma current and stored energy for five consecutive high performance shots, with statistically significant improvements relative to baseline, although we encourage caution in interpreting the statistics of the result due to the small sample size. As a test of the viability of this approach for performing small extrapolations in an incremental high performance campaign, which upcoming tokamaks like SPARC and ITER will undergo, we design a rampdown trajectory and perform a predict-first experiment by increasing the plasma current by 20%, from 140kA to 170kA, for a high $\beta_{N}$ plasma near the density limit, a scenario for which zero shots of rampdown data exists for TCV. In this extrapolation test, we a priori predict the dynamics of key quantities to within sufficient accuracy to successfully terminate the plasma on both attempts.
@@
Results
Experiment Overview
@@ Results
The reported experiment was conducted as a part of the 2024 TCV integrated control, high performance experimental campaign. Flattop plasmas operated at a high performance of $\beta_{N}>2.0$ and near the density limit with a highly elongated diverted geometry with $\kappa\approx 1.6$ and $q_{95}\approx 4$. Initial shots in the experiment operated at $I_{p}=140$ kA, henceforth known as the baseline high performance (HP) scenario, with a final extrapolation test at $I_{p}=170$ kA. Successful rampdowns from these scenarios require careful management of multiple plasma instability limits that can be exacerbated by details of the plasma trajectories. At present, a comprehensive understanding of disruptive limits remains an open problem, motivating many works on machine-learning based prediction of disruptions [35, 36, 37, 38, 39]. However, prior rampdown studies [8, 13, 9], a survey of disruption causes at TCV [40], and the present physics-based understanding of this high density scenario motivated constraints on four plasma parameters. Namely, we impose constraints on the Greenwald fraction, $f_{GW}$, the vertical instability growth rate $\gamma_{vgr}$, the plasma poloidal beta $\beta_{p}$, and the edge safety factor $q_{95}$. Managing the density limit, which is correlated with the Greenwald fraction, in this scenario is a particular challenge for fast terminations, as the relatively long particle confinement time-scale is a major constraint on the speed of the rampdown. The considered constraints are further discussed in the context of the reward function in the Methods.
@@
To address the problem, a NSSM dynamics model was trained on a modest dataset of past rampdowns, which contains only 5 incomplete rampdowns in the relevant high performance parameter space, as shown in Figure 2A-B. This NSSM is then used in a reinforcement learning (RL) environment to optimize a reward function, designed to minimize time to a goal plasma current of 404040kA and stored energy of 0.50.50.5kJ without disrupting, as shown in Figure 2C. The action space was chosen to be plasma current, $I_{p}$, shaping parameters $\kappa$ and $a_{minor}$ and neutral beam injection (NBI) power $P_{NBI}$. User-specified constraints were set on the Greenwald fraction $f_{GW}$, safety factor $q_{95}$, vertical instability growth rate $\gamma_{vgr}$ as calculated with the method in [41], and poloidal beta $\beta_{p}$. The optimized action trajectories were then manually programmed into the TCV plasma control system (PCS). The details of the reward function, chosen limits, and PCS programming process are further discussed in the Methods.
Statistical significance of control results in fusion is typically difficult to establish due to the scarcity of experimental time and relevant data-points. This experiment also faces this challenge, given the rampdown experiment involved only nine shots, two of which were dedicated to debugging a legacy software issue, with only five rampdowns in the database near the relevant high performance regime with $\beta_{N}>2$. We use these five shots as our control set and define two test sets: one with the debugging shots and one without. As shown in Figure 2D, the Mann-Whitney U test [42] shows a statistically significant improvement in $W_{tot}$ , ($p<0.05$), at time of plasma termination of the experimental rampdowns for both definitions of the test set. Improvements in $I_{p}$ are only statistically significant when we do not include the debugging shots. While the results of this statistical test are encouraging, we urge caution in its interpretation given the small sample sizes involved, and the fact that tokamaks are highly drifting distributions in practice, with uncontrolled variables such as wall conditioning making a meaningful impact on experimental outcomes.
@@
Every shot involved in the experiment is shown in Figure 3, showing improvements on $I_{p}$ and $W_{tot}$ at time of plasma termination over the course of the experimental runs. The un-optimized baseline rampdown trajectory for this scenario disrupted at relatively high current and stored energy in #81101 and #81102 at $I_{p}\approx 80$kA and $W_{tot}\approx 4$kJ. The experiment proceeded iteratively, with re-training of the NSSM on new data and trajectories done after shots #81635, #81745, #81751, and #81830. A preliminary optimized trajectory was deployed in TCV #81635, which reached the goal $I_{p}$ and $W_{tot}$ before disrupting, but post shot analysis showed poor radial control and tracking of the target shape, which was determined to be due to a legacy software issue detailed in Figure S1 in the Supplementary Information. Shots #81741 and #81745 were spent resolving this issue, with it properly resolved in #81751, as shown in Figure S1. #81751 still disrupted due to a VDE, which was found to be due to a large sensitivity of $\gamma_{vgr}$ to small control errors in the inner gap. After #81751, an uncertainty distribution on gap errors was added to the RL training environment to gain robustness to this uncertainty, with subsequent shots experiencing similar control errors but without similar increases in $\gamma_{vgr}$, demonstrating the importance of designing trajectories with robustness to real-world uncertainties, as further discussed in Section Preventing VDEs by Designing for Robustness to Control Error.
@@
Medium-Scale Validation of NSSM Predictions
@@
The model underwent two training phases: an initial training phase on a wider dataset with 311 shots in the training dataset and 131 shots in the validation dataset. To improve the predictive power of the model for the relevant scenario, we fine-tune just the confinement scaling of the model by training only on 44 shots with $I_{p}\leq 200$kA, with all other model weights frozen. Due to the relatively small size of the fine-tuning dataset, we did not separate out a validation or testing dataset for this fine-tuning phase. As shown in Figure 4, the trained model is able to predict the time-dependent dynamics of key 0D kinetic and disruptive quantities to within tens of percent for full rampdowns in the validation dataset, even in the 95th percentile of error. The percent errors for $\gamma_{vgr}$ can be relatively large, but, as shown in Figure S3 in the Supplementary Information, this is largely attributable to the small value of $\gamma_{vgr}$ of limited plasmas as the absolute error is relatively low.
@@
The NSSM was initially developed with a neural network predictor for the kinetic profiles on the full $\rho$ grid, and initial training runs found that the profile predictor can accurately predict kinetic profiles given the set of 0D scalars specified in Table 1. Figure 5 provides an example comparison of predictions of the $T_{e}$ and $n_{e}$ profiles against Thomson measurements for a full shot in the validation dataset, showing accurate prediction across all phases of the shot. This result corroborates previous findings at NSTX-U that neural networks can accurately predict kinetic profiles given a set of similar 0D scalars [43]. Given this result suggests most of the relevant profile information is implicitly captured by 0D scalars, the profile predictor was disabled prior to running experiments to accelerate training, hence reported predictions of kinetic profiles are not predict-first. This result also suggests a structured data-driven approach to modeling tokamak transport merits further research, in parallel with several ongoing principles-based efforts [44, 45, 46]. Another noteworthy feature of this profile predictor that should be explored in future work is its ability to function as a Thomson up-sampler, as the input variables are all sampled at a higher time resolution, 1ms, than the TCV Thomson Scattering system, which takes measurements every 17ms.
Preventing VDEs by Designing for Robustness to Control Error
The fact that the trajectory in #82875 is more robust to control error than in #81751 is corroborated by an analysis using the physics-based model for $\gamma_{vgr}$ used in this work [41]. Minor radius variations were introduced to the RL designed equilibrium trajectories for the two shots, yielding distributions of $\gamma_{vgr}$ trajectories. Figure 6 shows the conclusion that the trajectories in #82875 almost uniformly have lower $\gamma_{vgr}$ than #81751, and stay largely within the soft constraint specified in the RL training process, with the exception of the initial phase of the rampdown process as the flattop equilibrium has a large $\gamma_{vgr}$.
Improving Terminations by Incremental Re-training
Both the NSSM and trajectories were incrementally re-trained on newly generated data from experimental run-days, which resulted in more robust and faster rampdowns for the baseline high performance scenario, as shown in Figure 7. The speed of the model enabled re-training of both the model and trajectories in fewer than ten hours total on a single A100 GPU. The un-optimized solution in #81101 involved a NBI power rampdown while keeping constant plasma current, to allow for a decrease in density to avoid the Greenwald limit, a solution which the RL approach initially decided on as well, as shown with #81751, with an even more conservative current ramp and introducing a reduction in $\kappa$. As discussed in Preventing VDEs by Designing for Robustness to Control Error, this shot resulted in a VDE, and, with the introduction of an uncertainty distribution on $g_{HFS}$, the solution in #81830 resulted in less of a reduction in the minor radius $a_{minor}$, which helped eliminate the $\gamma_{vgr}$ spikes. Subsequent dynamics model training and trajectory optimization resulted in a solution in #82876 which allows for an immediate reduction in $I_{p}$ without running into a density limit, highlighting the ability for the workflow to assist in gradually making improvements. All optimized trajectories involved a fast initial drop in $P_{NBI}$, followed by a slower ramp phase, although the rates and critical points for the transition differed from shot to shot.
Predict-First Results for the Extrapolation Test
As shown in Figure 8, experimental results from #82878 largely fell within this distribution, with accurate predictions of the stored energy and density dynamics. Arguably the largest sources of error came from unreliable control of the plasma shape, contributing to errors in quantities like the rotational transform $\iota_{95}\equiv\frac{1}{q_{95}}$, and also leading to an earlier than expected HL back-transition. #82878 also started near the density limit, a challenging situation which motivated the introduction of a delay to the $I_{p}$ ramp in the baseline scenario, but the RL algorithm was able to determine a trajectory to immediately decrease $I_{p}$, which is desirable, while keeping $f_{GW}$ roughly constant. #82877 fell further out of distribution due to a loss of NBI power, and the presence of a neo-classical tearing mode (NTM) at the beginning of rampdown that did not exist in TCV #82878, as shown by Figure S5 in the Supplementary Information. Fortunately, these un-modeled off-normal events (ONEs) did not take the plasma far enough out of distribution to cause a disruption. As discussed earlier, the profile predictor was removed to help accelerate trajectory optimization, but post shot evaluation of the profile predictor on the 0D scalars generated by the training environment, shown in Figure 8, shows reasonable agreement against experimental Thomson measurements.
Discussion
[…]
The developed RL approach for designing robust trajectories yielded promising improvements in the plasma current and stored energy at time of termination, with incremental re-training improving the ramp speed. This result represents one of the first successful demonstrations of trajectory design with robustness to real-world uncertainties for tokamaks, which has historically been infeasible due to the computational cost of simulation. A degree of statistical significance is shown, but the sample size is still relatively small; a larger scale study would more thoroughly determine the efficacy of the approach. Although a large set of uncertainties was accounted for, detailed further in Table S1 in the Supplementary Information, experimental results involved additional uncertainties, such as the NBI failure in #82877, that still need to be addressed to further improve the robustness of trajectories. Robustness to hardware failure is of particular interest for future work as an exhaustive survey of disruption causes at JET has revealed hardware failure as a significant contributing factor to disruptions [48]. It is also noteworthy that the RL designed action trajectories tended to be relatively simple, suggesting that the key important ingredient is the fast and parallelized simulation model, as a human expert may be able to find similar trajectories if given access to the simulation model.
Methods
The Neural State-Space Model
$\dot{\mathbf{x}}$ $=A\mathbf{x}+B\mathbf{a}$ (1a)
$\mathbf{o}$ $=C\mathbf{x}+D\mathbf{a}$ (1b)
$\dot{\mathbf{x}}=NN_{\theta}(\mathbf{x})$ (2)
$\dot{\mathbf{x}}(t)$ $=f_{\theta}(\mathbf{x},\mathbf{a})$ (3a)
$\mathbf{o}(t)$ $=O_{\theta}(\mathbf{x},\mathbf{a})$ (3b)
The Dynamics Function $f_{\theta}(\mathbf{x},\mathbf{a})$
We begin by defining the following confinement laws:
$\text{tanhHeaviside}(x)$ $\equiv\frac{1}{2}(\tanh(kx+1))$ (4e)
where the parameters to be learned include all coefficients $c_{*}$ and neural network parameters. The laws are structured to multiply a portion corresponding to L-mode, a neural network correction factor, and an H-mode correction factor. The L-mode term reflects standard confinement scalings, but with the introduction of a $\dot{I}_{p}$, which was found to help better capture short-term effects of ramping plasma current. The neural network output includes a tanhClip final activation that constrains it’s output to the range [0.75,1.25][0.75,1.25][ 0.75 , 1.25 ], thus controlling the maximum adjustment the network is allowed to make. The hmode function includes a tanhHeaviside function which provides a smooth transition between one to zero once $P_{input}$ falls below the learned back-transition threshold, which is structured to reflect the Martin scaling [52]. Note that the use of the hmode function output as a power de-activates the H-mode correction term once hmode transitions from one to zero. While, in principle, the neural network should be able to learn the effects of H-mode implicitly, we found that adding an explicit H-mode correction factor helped improve model predictions in our low-data regime. The $k$ factor controls the smoothness of both the tanhClip and tanhHeaviside functions.
where $\sigma$ is the sigmoid function, $NN_{ohm,rad}$ is a network that predicts two quantities; the first is multiplied by $I_{p}^{2}$ to serve as an Ohmic heating term, and the second is multiplied by density and volume to serve as the radiated power term. $NN_{wall}$ is included to account for possible wall fueling effects when in a limited configuration, and is multiplied by an exponential in the HFS gap to de-activate it when diverted. Additional simple constants are included to account for fueling from both NBI and gas puffing. We note that, in both cases, the included terms do not capture important state dependencies and time delays, but they proved sufficient for this use case. The dynamics of density times volume are predicted; in cases where density itself is used (e.g. to compute the Greenwald Fraction), the following volume approximation is used to recover density:
The Observation Function $O_{\theta}(\mathbf{x},\mathbf{a})$
where $\beta_{p}$ is computed in accordance to the LIUQE definition [53], $f_{GW}$ is the usual Greenwald Fraction [10], $q_{95}$ is the approximation given in [54] with the squareness factor set to 1, $NN_{vgr}$ is a multi-layer perceptron (MLP with GELU activation and a scaled sigmoid output, and $NN_{prof}$ is a neural-operator based profile predictor, discussed further in the next subsection.
Neural Operator Based Profile Predictor
$f_{out}=\mathcal{F}_{\theta}(f_{in})$ (8)
And we predict these profiles using a set of 0D scalars, where every scalar is a control parameter except stored energy $W_{tot}$ and $\bar{n}_{e,20}$. The full set of input and output parameters is specified in Table 1. During training, the $\rho$ grid corresponding to the dataset is chosen to evaluate the basis functions, but arbitrary alternative grids can be used during inference time.
Training Methods
Training of the NSSM involved two stages. First, $NN_{ohm,rad}$, $NN_{vgr}$, and $NN_{prof}$ are trained independently of the rest of the model on time-independent samples to predict their respective quantities. These “pretrained” models are then integrated into the NSSM, where they are further trained jointly with the rest of the model through the time-dependent process specified in Figure 9. The AdamW optimizer [57] with an exponential decay learning rate schedule is used for every training run . All NNs used in the dynamics function $f$ and $NN_{vgr}$ are simple multi-layer perceptrons with GELU [58] activations on the hidden state and tanhClip functions as final activations to constrain their outputs to reasonable ranges. The profile predictor is further detailed earlier in the methods. Hyperparameters for the optimizer and model sizes are optimized via Bayesian Optimization using the method implemented in the Weights and Biases platform [59], which was used in this work for experiment tracking. The final set of hyperparameters are detailed in Table S2 in the Supplementary Information.
Training Data Distribution
The dataset used for training models in this work consists of the 442 most recent shots with rampdowns that are at least partially complete and sufficient diagnostic availability, gathered with DEFUSE (Disruption and Event analysis framework for FUSion Experiments) [60]. The initial training phase involved training on 311 shots of data, with the rest of the dataset used for validation. After the initial training phase, the model is further trained on a fine-tuning dataset of 44 shots, during this phase all of the model weights except those in the $\tau_{E}$ and $\tau_{N}$ hybrid confinement laws described in The Dynamics Function $f_{\theta}(\mathbf{x},\mathbf{a})$ are frozen. As shown in Figure S8 in the Supplementary Information, the dataset consists of only five shots of data anywhere near the relevant high-performance region.
Reward Function
One challenge with specifying constraint limits is the difference in severity of violating different constraints, and the, at times, weak correlations between physical quantities and disruptions. To address this issue, we partition constraints into “soft” constraints, which incur a small penalty to discourage, but not forbid, the algorithm from finding solutions that violates these limits, and “hard” constraints which incur a large penalty to strictly enforce constraint violation. We note that while methods in the constrained optimization literature often mathematically express constraints separately from the objective function being optimized, most practical implementations of constrained optimization algorithms enforce constraints by rewriting constraints as penalty terms in the objective function [62, 63], an approach we also adopt. Stochastic optimization across a distribution of outcomes introduces a challenge: trying to avoid limits for every scenario will likely result in excessively slow and conservative solutions [64], which itself poses it’s own risk. To address this challenge, we utilize chance-constraints, a technique often utilized in the autonomous driving literature [65, 66], and only activate the constraint if violation probability exceeds a certain threshold. In this set of experiments, this threshold is chosen as 5%. Reward function parameters used for the final four shots are shown in Table S2 in the Supplementary Information.
Uncertainty Model
In experimental reality, the time evolution of plasma dynamics is highly nonlinear and subject to considerable amounts of uncertainty, as evidenced by the two same-scenario shots shown in Figure 8 which begin at drastically different initial conditions. To design trajectories that have robustness to large variability and off-normal events, we defined an uncertainty model for the RL training environments, and sampled from this uncertainty model for each training environment used during training. The uncertainty model includes random variables for both the initial state of the plasma during rampdown and disturbances/model uncertainties that affect the time-varying plasma dynamics. To account for the fact that accidental H-L back-transition implies the initial state of the plasma may start in either H or L-mode, the initial state distribution is modeled as a bi-modal mixture model, with a 50% chance of any given RL training environment starting in either H or L-mode. In some cases, uncertainty distributions could easily be quantified from past experimental data (such as tracking error in the plasma current), or from model prediction accuracy (such as $\gamma_{vgr}$), but in other cases the distribution was chosen in ad-hoc fashion, upon identifying additional sources of uncertainty in experiment. Table S1 in the Supplementary Information summarizes the random variables, parameterized distributions, and quantification methods used in this work. As discussed in Section Predict-First Results for the Extrapolation Test, this uncertainty model proved to be non-exhaustive in experiment. In addition, the uncertainty model employed does not account for time-varying fluctuations in uncertain variables; future work should employ time-varying stochastic processes. Both of these limitations further highlight the need to advance experimental uncertainty quantification and robust control in the context of fusion plasma control.
RL Methods
$\mathbf{a}=\pi(\mathbf{o})$ (11)
$\mathbf{a}=\pi(t)$ (12)
Given that time is the only observable, but there exists different physical conditions in the parallel training environments which are unobservable to the policy, the reward maximization process yields a trajectory that is designed to succeed across the different conditions specified in Subsection Uncertainty Model. After an initial trial with Proximal Policy Optimization (PPO) [67], we found OpenAI-ES, an evolutionary strategy (ES) designed for policy optimization, to work better in practice [68]. This is possibly explained by the theoretical analysis given in the paper introducing OpenAI-ES, which suggests that RL problems with long time horizons and actions that have long-lasting effects may be better solved with ES approaches than the dominant paradigm of policy gradient methods [68]. The policy $\pi$ was parameterized by a multi-layer perceptron with two hidden layers of width 64 and used relu activations with a hyperbolic tangent final activation to constrain the action space. Hyperparameter sweep of the architecture was not employed and would be worthwhile to investigate for future works.
Deployment to TCV
Data Availability
Data to generate figures found in this paper are available in the source code repository at https://doi.org/10.5281/zenodo.16621120. The complete dataset used for training and validation can be obtained by contacting Alessandro Pau or Allen M. Wang.
Code Availability
Source code is provided at https://doi.org/10.5281/zenodo.16621120.
Acknowledgments
Allen M. Wang and Cristina Rea were funded in part by Commonwealth Fusion Systems. This work has been carried out within the frame-work of the EUROfusion Consortium, via the Euratom Research and Training Programme (Grant Agreement No 101052200 - EUROfusion) and funded by the Swiss State Secretariat for Education, Research, and Innovation (SERI). Alessandro Pau, Olivier Sauter, Anna Vu, Cristian Galperti, Antoine Merle, Yoeri Poels, Cristina Venturini, Federico Felici, and Stefano Marchioni received support from EUROfusion and SERI. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission, or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them. The authors would like to acknowledge the Engaging cluster, managed by the MIT Office of Research Computing and Data, which was used in this work for model training.
Author Contributions Statement
Allen M. Wang led the project, developed the dynamics modeling approach and RL problem formulation, and led writing of the paper. Alessandro Pau developed the machine learning dataset, high performance integrated control scenario, and integration of the approach into TCV. Oswin So and Charles Dawson worked with Allen M. Wang to develop the RL training environments and methods. Olivier Sauter enabled TCV integration, and identified and debugged the radial observer issue. Cristina Rea managed the collaboration, advised the project on disruptions and machine learning, and contributed to manuscript drafting and revisions. Mark D. Boyer advised the project from a controls perspective and motivated the approach taken for dynamics modeling. Anna Vu and Cristian Galperti enabled integration with the integrated control system. Chuchu Fan advised Allen M. Wang, Oswin So, and Charles Dawson on controls and RL approaches. Antoine Merle developed and assisted with MHD equilibrium codes. Yoeri Poels and Cristina Venturini helped with the TCV dataset. Federico Felici developed equilibrium and plasma control infrastructure that enabled this work and advised development of the growth rate calculation method. Stefano Marchioni developed the growth rate calculation method used in this work.
Competing Interests Statement
Mark Dan Boyer is an employee of Commonwealth Fusion Systems. Allen M. Wang and Cristina Rea are funded in part by Commonwealth Fusion Systems.
Tables